OpenStudy (anonymous):

find the exact value of cos(sin^-1(3/5))

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

I got \( \huge \frac{4}{5} \)

OpenStudy (michele_laino):

If \theta is such that: \[\Large \theta = \arcsin \left( {\frac{3}{5}} \right)\] then you have to compute this: \[\Large \cos \theta \]

OpenStudy (anonymous):

were not suppose to use calculator :/

OpenStudy (anonymous):

whats arc sin?

OpenStudy (michele_laino):

ok! now we have: \[\Large \sin \theta = \frac{3}{5}\] am I right?

OpenStudy (michele_laino):

it is another way to mean sin^-1

OpenStudy (michele_laino):

so, if we apply the fundamental identity: \[\Large {\left( {\cos \theta } \right)^2} + {\left( {\sin \theta } \right)^2} = 1\] we can write: \[\Large \cos \theta = \pm \sqrt {1 - {{\left( {\sin \theta } \right)}^2}} = \pm \sqrt {1 - {{\left( {\frac{3}{5}} \right)}^2}} = ...?\]

OpenStudy (anonymous):

let 2 sin^-1(3/5) = A sin^-1(3/5) = A/2 sin(A/2) = 3/5 cos A = 1 - 2sin^2(A/2) = 1 - 2(9/25) = 7/25 Thus cos [ 2sin^-1(3/5)] = cos A = 7/25 I found this answer but where could they have gotten 9 from?

OpenStudy (michele_laino):

no, we have: cos(A/2) =sqrt(1 - sin^2(A/2)

OpenStudy (anonymous):

wait so the way they did it is wrong?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (michele_laino):

it is not necessary to work with A/2. Please apply my procedure

OpenStudy (michele_laino):

I used +/- since we have two square roots

OpenStudy (michele_laino):

for example: \[\Large \sqrt 4 = \pm 2\]

OpenStudy (anonymous):

Wait so how would your work look like then?

OpenStudy (anonymous):

Because your procedure is confusing me even more lol

OpenStudy (michele_laino):

since we have: \[\Large \sin \theta = \frac{3}{5}\] then we can write: \[\Large \cos \theta = \pm \sqrt {1 - {{\left( {\sin \theta } \right)}^2}} = \pm \sqrt {1 - {{\left( {\frac{3}{5}} \right)}^2}} = ...?\]

OpenStudy (anonymous):

is this a half angle formula?

OpenStudy (michele_laino):

no, it is not a bisection formula

OpenStudy (michele_laino):

hint: \[\Large \sqrt {1 - {{\left( {\frac{3}{5}} \right)}^2}} = \sqrt {1 - \frac{9}{{25}}} = \sqrt {\frac{{25 - 9}}{{25}}} = ...?\]

OpenStudy (anonymous):

Is there another way to do it? becauase our teacher didn't show us that formula.

OpenStudy (michele_laino):

Sincerely speaking, it is the unique way that I know

OpenStudy (anonymous):

but whats wrong with the method i used? im still not understanding

OpenStudy (michele_laino):

your method is correct, nevertheless you have to find cos(A/2) not cos(A)

OpenStudy (michele_laino):

you have used the subsequent bisection formula: \[\Large {\left\{ {\sin \left( {\frac{\theta }{2}} \right)} \right\}^2} = \frac{{1 - \cos \theta }}{2}\]

OpenStudy (anonymous):

This question just got worse so I used something else but idk if its riight..

OpenStudy (michele_laino):

the computation of your image is correct

OpenStudy (anonymous):

So is 7/25 right then??

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we can write this: \[\Large \sin \theta = \frac{3}{5} \Rightarrow \cos \left( {2\theta } \right) = \frac{7}{{25}}\]

OpenStudy (anonymous):

@Michele_Laino lol what does that mean tho? XD can you explain it?

OpenStudy (michele_laino):

it is simple: since \[\Large \sin \theta = \frac{3}{5}\] then: \[\Large \cos \theta = \pm \sqrt {1 - {{\left( {\sin \theta } \right)}^2}} = \pm \sqrt {1 - {{\left( {\frac{3}{5}} \right)}^2}} = \pm \frac{4}{5}\] Now we use the duplication formula: \[\Large \begin{gathered} \cos \left( {2\theta } \right) = {\left( {\cos \theta } \right)^2} - {\left( {\sin \theta } \right)^2} = \hfill \\ \hfill \\ = {\left( {\frac{4}{5}} \right)^2} - {\left( {\frac{3}{5}} \right)^2} = \frac{{16}}{{25}} - \frac{9}{{25}} = \frac{{16 - 9}}{{25}} = \frac{7}{{25}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ohh alright its the double angle formula too!